OpenStudy (anonymous):

What is the similarity ratio of a cube with volume 1,728 m3 to a cube with volume 19,683 m3?

OpenStudy (anonymous):

9:4 4:9 144:729 729:144

OpenStudy (ranga):

I suppose they are asking for the ratio of length of the side of cube1 to the length of the side of cube 2. Take the cube root of each number. The ratio will be: cube root of 1728 : cube root of 19,683. Simplify the ratio.

OpenStudy (anonymous):

12:27?

OpenStudy (ranga):

divide both by 3 to simplify.

OpenStudy (anonymous):

4:9

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

think you could help me on one more with ratios ?

OpenStudy (ranga):

I can try.

OpenStudy (anonymous):

The measures of the angles of a triangle are in the extended ratio 4:12:14. What is the measure of the smallest angle? 6° 24° 72° 84°

OpenStudy (ranga):

The three angles of a triangle add to 180 degrees. The smallest angle will be 4 / (4+12+14) of 180 degrees.

OpenStudy (anonymous):

hmmm i dont really understand that :/ if you do 4/(4+12+14) it gets like .133333333

OpenStudy (ranga):

Multiply by 180.

OpenStudy (ranga):

A better way to do that is: 4/(4+12+14) x 180 = 4/30 x 180 = 4 x 6 = 24

OpenStudy (anonymous):

i see. when i multiplied by 180 i got another odd number

OpenStudy (anonymous):

but thank you! I really needed that help

OpenStudy (ranga):

you are welcome.